# Online Decision-Focused Learning

### v2

@@ 2 Framework
From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ using a variant of either the Follow-The-Leader algorithm [FTL, 45] or Online Gradient Descent [OGD, 81], applied to the objective function $f_{t}$ defined in (3), thus specifying the choice of $\mathsf{Alg}_{t}$. However, these algorithms are designed for single-level optimization problems and require access to (sub-)gradients of the objective. In our setting, this requirement is problematic because the function $f_{t}$ does not admit an informative gradient: by construction, computing the gradient of $f_{t}$ involves differentiating the mapping $\theta\mapsto w_{t}^{\star}(\theta)$, which is generally not (sub-)differentiable. Indeed, $w^{\star}_{t}$ minimizes a linear function over the convex polytope $\mathcal{W}$, thus $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$.
To address this issue, following [74], we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,
@@
which is a regularized approximation of $w^{\star}_{t}$. When this surrogate is continuously differentiable—which holds for our choices of $\mathcal{R}$ in Section˜3—we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine. This approach amounts to minimizing the surrogate
@@
It remains to determine what regularizer $\mathcal{R}$ to choose in (6). We recall that our aim is to obtain a $\tilde{w}_{t}(\theta)$ in (5) which is differentiable for any $\theta\in\Theta$. A possible strategy is to choose $\mathcal{R}$ so $\tilde{w}_{t}$ remains in the strict interior of $\mathcal{W}$. In this case, $\tilde{w}_{t}(\theta)$ is differentiable in a neighborhood of any $\theta\in\Theta$ and admits a close-form Jacobian $\nabla\tilde{w}_{t}(\theta)$ by the implicit function theorem. A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function:
@@
The bi-level structure of the problem yields unexpected properties. In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26)—which the case with both log-barriers and negative entropy—and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ may not be convex, but is lipschitz (see Lemmas˜5 and 1). This prevents us from directly using known online convex optimization algorithms [37]. However, a recent line of research [3, 67, 75] has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous—which is the case with log-barriers or negative entropy (see Lemmas˜1 and 5). These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret. In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli [67].
@@
An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
@@
The notion of an approximate oracle formalizes that, in non-convex settings, we cannot rely on subroutines that provably attain a global minimizer of $\tilde{f}_{t}$ at time $t$ (as offline gradient descent would in convex problems). Instead, we must settle for local minimizers, whose quality is characterized by a parameter $\xi$—which vanishes in favorable loss landscapes. As a concrete example, consider the stochastic gradient descent (SGD) algorithm as $\mathbf{O}_{\xi}$. A large body of work has shown that, even when $\Theta$ is a class of deep neural networks, SGD can converge to local minimizers [31, 55, 58, 20], justifying that $\mathbf{O}_{\xi}$ is indeed an approximate oracle. A more detailed discussion is provided in Appendix˜B.
We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see [38, 6, 80, 36]. However, such a definition is not meaningful in the DFL framework, where gradients are zero or undefined due to the structure of the problem.
@@ 3.1 Decision-Focused Follow The Perturbed Leader
Our first algorithm is displayed in Algorithm˜1, where $\mathrm{Exp}(\eta)$ refers to an exponential distribution with parameter $\eta>0$. It is inspired from the classic Follow-the-Leader approach [45], which consists in making a decision minimizing the sum of objective functions observed so far. When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader [43, 67]. We employ this strategy as $\mathsf{Alg}_{t}$ to update $\theta_{t+1}$ in Algorithm˜1. Note that the oracle in line 6 of Algorithm 1 crucially works with the regularized losses $\tilde{f}_{1},\ldots,\tilde{f}_{t}$. This allows to use in practice gradient-based methods to obtain an approximate minimizer as per our discussion in Section˜3, since these surrogate losses are differentiable and Lipschitz.
@@
Note that Algorithm˜1 requires determining $w^{\star}$, which can be efficiently achieved by classic numerical solvers in reasonable settings. However, we show in Appendix˜F that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
We provide a theoretical guarantee on the convergence of Algorithm˜1 in the form of the following static regret bound.
@@
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}\sqrt{n}T^{-1/4}+\xi}}\right).$
@@
It is informative to compare our guarantee to existing bounds in the literature. [67], who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension. Their faster rate in $T$ comes from the fact that they tackle a simpler, single-level problem as compared to our bi-level, non differentiable setting. Indeed, we need to regularize the inner problem to overcome non-differentiability as discussed in Section˜3. This introduces an additional trade-off on the regularization strength $\alpha$ on top of the usual trade-off in the learning rate $\eta$, which is reflected in our rate.
@@ 3.2 Decision-Focused Online Gradient Descent
While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in [67] are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another. To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent [81]. This procedure, which we call DF-OGD, is presented in Algorithm˜2.
In words, at each time $t\in[T]$, Algorithm˜2 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$ and computes one of its near minimizer via the offline oracle from Definition˜1. Next, to handle non-convexity, it determines the gradient $\nabla\tilde{w}_{t}(u_{t})$, evaluated in a point $u_{t}$ drawn uniformly at random in $\theta_{t},\vartheta_{t}]$. From there, it updates $\theta_{t+1}$ through the classical gradient step scheme of OGD.
@@
Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$. Then, using $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
@@
The complete statement of the main bound, including constants is provided in Section˜G.3. Our dynamic regret bound naturally depends on $P_{T}$, which captures the problem’s variability by measuring the cumulative distance between approximate minimizers over time. When $\xi=\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$, the average dynamic regret decreases at the rate $\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$. Notably, the bound is independent of the dimension of $\Theta$, and depends only mildly on the dimension of $\mathcal{W}$ through a $\ln\ln(d)$ factor. This constitutes a key strength of the optimistic strategy underlying DF-OGD, making it particularly well-suited for high-dimensional settings.
It is instructive to compare our guarantee with those established in recent studies on dynamic regret. For instance, [77] derive a $\mathcal{O}(\sqrt{T(P_{T}+1)})$ bound. However, their setting is not directly comparable to ours, as they consider a simpler single-level problem with differentiable and convex objectives. In contrast, our framework involves non-convex, non-differentiable losses due to the bi-level nature of decision-focused learning. More recently, [41] obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$. Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of $f_{t}$s. By comparison, our losses are neither Lipschitz (indeed, they may even be discontinuous due to the linearity of the lower-level problem) nor required to be stationary over time—the only restriction being H˜2 to hold. We view the ability of our algorithm to achieve efficient convergence despite these demanding conditions as a core contribution of our work.
@@ 4 Experiments
In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO [25].
@@
Our experimental setup is inspired by the knapsack example from Mandi et al. [54]. More precisely, we consider a decision maker who must pick at each $t\in[T]$ an object $v_{t}\in\mathcal{V}$ among $K$ items denoted $\mathcal{V}=\{1,2,\ldots,K\}$ with respective costs $\bar{g}_{t}(X)=(\bar{g}_{t,1}(X),\ldots,\bar{g}_{t,K}(X))\in[0,1]^{K}$ depending on some covariates $X\in\mathcal{X}$. At the beginning of each period, the decision-maker only observe covariates $X_{t}\in\mathcal{X}\subset\mathbb{R}^{p}$, and have at their disposal a parametric model $g:\Theta\times\mathcal{X}\rightarrow[0,1]^{K}$ to predict $\bar{g}_{t}$. Given their current parameter $\theta_{t}\in\Theta$, they predict item costs as $g(\theta_{t},X_{t})=(g_{1}(\theta_{t},X_{t}),\ldots,g_{K}(\theta_{t},X_{t}))$ and pick an item $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}g_{i}(\theta_{t},X_{t})$. This setting is depicted in Figure˜2 for $K=2$ in Appendix˜E. After having made their decision, the decision-maker observes the true item costs $\bar{g}_{t}(X_{t})$, and update $\theta_{t+1}$ for the next round based on this feedback. Note that this setting can directly be mapped in the simplex example of Remark˜1, since $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{\left\langle w,g(\theta_{t},X_{t})}}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$. For this reason, we instantiate our algorithms DF-OGD and DF-FTPL with the negative entropy regularizer, following Remark˜1.
@@
In this setting, we compare the performances of DF-FTPL (Algorithm˜1) and DF-OGD (Algorithm˜2) to two benchmarks. First, we implement Prediction-Focused Online Gradient Descent (PF-OGD). This strategy consists in training in an online manner the model $g$ at each timestep so it minimizes the statistical loss $\ell^{\mathrm{mse}}:(v,X\theta)\mapsto\mathopen{}\mathclose{{\left\|v-X\theta}}\right\|^{2}$, irrespective of the downstream decision problem. Then, decisions are greedily made based on the predictions of the model. This approach is formally described in Algorithm˜3 in Appendix˜E. Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from [25]. This algorithm introduces a differentiable and convex loss which upper-bounds the actual decision-focused loss. Given its effectiveness and versatility, it is considered as a very important benchmark in the literature.

### v3

@@ 3 Online algorithms for decision-focused learning
From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ through an algorithm $\mathsf{Alg}_{t}$ being either a variant of either the Follow-The-Leader algorithm (FTL, Kalai and Vempala, 2005) or Online Gradient Descent (OGD, Zinkevich, 2003), applied to the objective function $f_{t}$ defined in (3). However, these algorithms are designed for single-level optimization problems and require access to (sub-)gradients of the objective. In our setting, this requirement is problematic because the function $f_{t}$ does not admit an informative gradient: by construction, computing the gradient of $f_{t}$ involves differentiating the mapping $\theta\mapsto w_{t}^{\star}(\theta)$, which is generally not (sub-)differentiable. Indeed, $w^{\star}_{t}$ minimizes a linear function over the convex polytope $\mathcal{W}$, thus $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$.
To address this issue, following Wilder et al. (2019), we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$ and some $\alpha_{t}>0$
@@
which is a regularized approximation of $w^{\star}_{t}$. When this surrogate is continuously differentiable, which is the case for our choices of $\mathcal{R}$ in Section˜3, we can use $\nabla\tilde{w}_{t}(\theta)$ in our algorithmic routine. This approach amounts to minimizing the surrogate
@@
It remains to determine what regularizer $\mathcal{R}$ to choose in (6). We recall that our aim is to obtain a $\tilde{w}_{t}(\theta)$ in (5) which is differentiable for any $\theta\in\Theta$. A possible strategy is to choose $\mathcal{R}$ so $\tilde{w}_{t}$ remains in the strict interior of $\mathcal{W}$. In this case, $\tilde{w}_{t}(\theta)$ is differentiable in a neighborhood of any $\theta\in\Theta$ and admits a close-form Jacobian $\nabla\tilde{w}_{t}(\theta)$ by the implicit function theorem. A natural choice to force $\tilde{w}_{t}$ to remains in the interior of $\mathcal{W}$ is the corresponding log-barrier function
@@
The bi-level structure of the problem yields unexpected properties. In particular, even when the regularizer $\mathcal{R}$ is strongly-convex in (26), which the case with both log-barriers and negative entropy, and the model $g$ is simple (such as $g(\theta,X)=X\theta$), $\tilde{f}_{t}:\theta\mapsto\mathopen{}\mathclose{{\left\langle\tilde{w}_{t}(\theta),\bar{g}_{t}(X_{t})}}\right\rangle$ is lipschitz (see Lemmas˜5 and 1) but not convex in general. This prevents us from directly using known online convex optimization algorithms (Hazan, 2023). However, a recent line of research (Agarwal et al., 2019; Suggala and Netrapalli, 2020; Xu and Zhang, 2024) has developed online algorithms in the non-convex case, provided the losses are Lipschitz-continuous. These studies combine near-optimal oracles with perturbation techniques to establish sublinear bounds on the expected regret. In line with this literature, we also assume to have access to an approximate offline optimization oracle, a notion that appeared in a slightly modified form in Suggala and Netrapalli (2020).
@@
An $\xi$-approximate offline optimization oracle (or approximate oracle), adapted to a class $\mathcal{C}$ is a mapping $\mathbf{O}_{\xi}$ taking a function $f\in\mathcal{C}$ such that $f:\Theta\rightarrow\mathbb{R}$, and outputting $\vartheta=\mathbf{O}_{\xi}(f)\in\Theta$ satisfying:
@@
This notion of an approximate oracle reflects the fact that, in non-convex settings, we cannot rely on subroutines that provably find a global minimizer of $\tilde{f}_{t}$ at time $t$ (as is possible in convex problems with offline gradient descent). Instead, we must be content with local minimizers. Their quality is measured by a parameter $\xi$, which becomes small in favorable loss landscapes. When $\mathcal{C}$ contains differentiable and Lipschitz functions, the stochastic gradient descent (SGD) algorithm can be used for the oracle $\mathbf{O}_{\xi}$. Indeed, a large body of work has shown that, even with models as complicated as deep neural networks, SGD can converge to local minimizers (Ghadimi and Lan, 2013; Mertikopoulos et al., 2020; Patel and Zhang, 2021; Cutkosky et al., 2023), justifying its use as an approximate oracle. A more detailed discussion is provided in Appendix˜B.
We conclude this remark by mentioning that it is possible to conduct a non-convex analysis without an approximate oracle by considering the weaker notion of local regret defined as $\sum_{t\in[T]}\mathopen{}\mathclose{{\left\|\nabla F_{t}(\theta_{t})}}\right\|$; see (Hazan et al., 2017; Aydore et al., 2019; Zhuang et al., 2020; Hallak et al., 2021). However, such a definition is not meaningful in the DFL framework, where gradients are zero or undefined due to the structure of the problem.
@@ 3.1 Decision-Focused Follow The Perturbed Leader
Our first algorithm is displayed in Algorithm˜1, where $\mathrm{Exp}(\eta)$ refers to an exponential distribution with parameter $\eta>0$. It is inspired from the classic Follow-the-Leader approach (Kalai and Vempala, 2005), which consists in making a decision minimizing the sum of objective functions observed so far. When losses are non-convex, it is common to inject random noise at each step to regularize the total cost function, resulting in an approach known as Follow-the-Perturbed-Leader (Hutter and Poland, 2005; Suggala and Netrapalli, 2020). We employ this strategy as $\mathsf{Alg}_{t}$ to update $\theta_{t+1}$ in Algorithm˜1. Note that the oracle in line 6 of Algorithm 1 crucially works with the regularized losses $\tilde{f}_{1},\ldots,\tilde{f}_{t}$. This allows to use in practice gradient-based methods to obtain an approximate minimizer as per our discussion in Section˜3, since these surrogate losses are differentiable and Lipschitz.
@@
Note that in line 3, the algorithm takes the best action $w^{\star}_{t}(\theta_{t})$ given the predicted cost $g(\theta_{t},X_{t})$ (the regularized action $\tilde{w}_{t}$ is only needed to compute $\tilde{f}_{t}$). In most practical settings, $w^{\star}_{t}$ can be computed efficiently with standard numerical solvers. However, we show in Appendix˜F that only being able to determine $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ such that $\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),\underline{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w^{\star}_{t}(\theta_{t})}}\right\rangle\leqslant\kappa$ for some $\kappa>0$ only shifts the regret bounds of the next section by $\kappa$.
We provide a theoretical guarantee on the convergence of Algorithm˜1 with the following static regret bound.
@@
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(m^{3/4}n^{3/2}T^{-1/4}+\xi}}\right).$
@@
It is informative to compare our guarantee to existing bounds in the literature. Suggala and Netrapalli (2020), who also study non-convex online learning, achieves a rate of $\mathcal{O}(T^{-1/2})$ as long as their offline oracle satisfies $\xi=\mathcal{O}(T^{-1/2})$, at the cost of a degraded $m^{3/2}$ dependency on the dimension. Their faster rate in $T$ comes from the fact that they tackle a simpler, single-level problem as compared to our bi-level, non differentiable setting. Indeed, we need to regularize the inner problem to overcome non-differentiability as discussed in Section˜3. This introduces an additional trade-off on the regularization strength $\alpha$ on top of the usual trade-off in the learning rate $\eta$, which is reflected in our rate.
@@ 3.2 Decision-Focused Online Gradient Descent
While DF-FTPL (Algorithm˜1) benefits from a converging static regret bound, the techniques involved in Suggala and Netrapalli (2020) are not enough to reach dynamic regret guarantees, which is particularly relevant in highly non-stationary environments where the optimal decisions may change significantly from one round to another. To this end, we go beyond the Follow-the-Leader approach and propose an original algorithm based on the celebrated Online Gradient Descent (Zinkevich, 2003). This procedure, which we call DF-OGD, is presented in Algorithm˜2.
In Algorithm˜2, at each time $t\in[T]$, a decision $w_{t}^{\star}(\theta_{t})\in\mathcal{W}$ is made based on the current parameter $\theta_{t}\in\Theta$. The feedback $\bar{g}_{t}(X_{t})\in\mathcal{Z}$ is then observed and used to define the surrogate objective $\tilde{f}_{t}$. Then, a near-minimizer of this objective is computed using the offline oracle from Definition˜1. To address non-convexity, the gradient $\nabla\tilde{w}t(u_{t})$ is evaluated at a point $u_{t}$ drawn uniformly at random from $\theta_{t},\vartheta_{t}]$. Finally, the parameter is updated to $\theta{t+1}$ using the standard gradient step of OGD.
@@
Furthermore, assume $(t^{-1}(1+P_{t}))_{t\geqslant 1}$ is non-increasing almost surely with $P_{t}=\sum_{s=1}^{t-1}\mathopen{}\mathclose{{\left\|\vartheta_{s+1}-\vartheta_{s}}}\right\|$. Then, using $\alpha_{t}\propto n^{-1/2}t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto n^{-1/2}t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to:
@@
Our dynamic regret bound naturally depends on $P_{T}$, which captures the problem’s variability by measuring the cumulative distance between approximate minimizers over time. When $\xi=\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$, the average dynamic regret decreases at the rate $\mathcal{O}((1+P_{T})^{1/4}T^{-1/4})$. Notably, the bound is independent of the dimension of $\Theta$, and depends only mildly on the dimension of $\mathcal{W}$ through a $\ln\ln(d)$ factor. This constitutes a key strength of the optimistic strategy underlying DF-OGD, making it particularly well-suited for high-dimensional settings.
It is instructive to compare our guarantee with those established in recent studies on dynamic regret. For instance, Zhang et al. (2018) derive a $\mathcal{O}(T^{-1/2}\sqrt{T(P_{T}+1)})$ bound. However, their setting is not directly comparable to ours, as they consider a simpler single-level problem with differentiable and convex objectives. In contrast, our framework involves non-convex, non-differentiable losses due to the bi-level nature of decision-focused learning. More recently, Huang and Wang (2025) obtained a $\mathcal{O}((1+P^{\infty}_{T})^{1/3}T^{-1/3})$ bound, where $P^{\infty}_{T}:=\sum_{t=1}^{T}\|f_{t+1}-f_{t}\|_{\infty}$. Yet, this rate is achieved under substantially more favorable conditions: a single-level problem with losses that are strongly convex or Lipschitz, an additional assumption of “quasi-stationary” and a more challenging path involving the full landscapes of the $f_{t}$s. By comparison, our losses are neither Lipschitz (indeed, they may even be discontinuous due to the linearity of the lower-level problem) nor required to be stationary over time—the only restriction being H˜2 to hold. We view the ability of our algorithm to achieve efficient convergence despite these demanding conditions as a core contribution of our work.
@@ 4 Experiments
In this section, we compare the performances of our algorithms DF-FTPL and DF-OGD to two important benchmarks, namely prediction-focused learning and SPO (Elmachtoub and Grigas, 2022).
@@
Our experimental setup is inspired by the knapsack example from Mandi et al. (2024). More precisely, we consider a decision maker who must pick at each $t\in[T]$ an object $v_{t}\in\mathcal{V}$ among $K$ items denoted $\mathcal{V}=\{1,2,\ldots,K\}$ with respective costs $\bar{g}_{t}(X)=(\bar{g}_{t,1}(X),\ldots,\bar{g}_{t,K}(X))\in[0,1]^{K}$ depending on some covariates $X\in\mathcal{X}$. At the beginning of each period, the decision-maker only observe covariates $X_{t}\in\mathcal{X}\subset\mathbb{R}^{p}$, and have at their disposal a parametric model $g:\Theta\times\mathcal{X}\rightarrow[0,1]^{K}$ to predict $\bar{g}_{t}$. Given their current parameter $\theta_{t}\in\Theta$, they predict item costs as $g(\theta_{t},X_{t})=(g_{1}(\theta_{t},X_{t}),\ldots,g_{K}(\theta_{t},X_{t}))$ and pick an item $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}g_{i}(\theta_{t},X_{t})$. This setting is depicted in Figure˜2 for $K=2$ in Appendix˜E. After having made their decision, the decision-maker observes the true item costs $\bar{g}_{t}(X_{t})$, and update $\theta_{t+1}$ for the next round based on this feedback. Note that this setting can directly be mapped in the simplex example of Remark˜1, since $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{\left\langle w,g(\theta_{t},X_{t})}}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$. For this reason, we instantiate our algorithms DF-OGD and DF-FTPL with the negative entropy regularizer, following Remark˜1.
@@
In this setting, we compare the performances of DF-FTPL (Algorithm˜1) and DF-OGD (Algorithm˜2) to two benchmarks. First, we implement Prediction-Focused Online Gradient Descent (PF-OGD). This strategy consists in training in an online manner the model $g$ at each timestep so it minimizes the statistical loss $\ell^{\mathrm{mse}}:(v,X\theta)\mapsto\mathopen{}\mathclose{{\left\|v-X\theta}}\right\|^{2}$, irrespective of the downstream decision problem. Then, decisions are greedily made based on the predictions of the model. This approach is formally described in Algorithm˜3 in Appendix˜E. Second, we compare our algorithms to an online version of the Smart Predict-then-Optimize (online SPO, see Algorithm˜4) approach from Elmachtoub and Grigas (2022). This algorithm introduces a differentiable and convex loss which upper-bounds the actual decision-focused loss. Given its effectiveness and versatility, it is considered as a very important benchmark in the literature.